# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v3

@@ 2.3 Benchmark and evaluation of Retrieval-Augmented Generation
- Domain | |
- Total Pages | |
- Avg.Tokens | |
- Avg.Data Type | |
@@
- Text | |
- Table | |
- Formula | |
- Chart | |
- Reading Order | |
@@
- String | |
- Numeric | |
- Yes/No | |
- List | |
@@
- Understanding | |
- Reasoning | |
- Single-page | |
- Multi-page |
@@ IV.6 Details in different domains
1. Question should represent a plausible inquiry that a person (who has not seen the page) might ask about the information uniquely presented on this page. The questions should not reference this specific page directly (by page number, pointing to a specific paragraph or figure, and never refer to the document using phrases like ’in the document’), nor should they quote the text verbatim. They should use natural language reflecting how someone might inquire about the page’s content without direct access. |
2. Question must contain all information and context/background necessary to answer without the document. Do not include phrases like "according to the document" in the question. |
3. Question must not contain any ambiguous references, such as ’he’, ’she’, ’it’, ’the report’, ’the paper’, and ’the document’. You MUST use their complete names. |
@@
- The question MUST be detailed and be based explicitly on information in the document. |
- The question MUST include at least one entity. |
- The context sentence the question is based on MUST include the name of the entity. For example, an unacceptable context is "He won a bronze medal in the 4 × 100 m relay". An acceptable context is "Nils Sandström was a Swedish sprinter who competed at the 1920 Summer Olympics." |
- The answer form should be as diverse as possible, including [Yes/No, Numeric, String, List]. |
- {additional_task_criteria} |
@@
- The answer to the combined question should be the same as the answer2. |
- It must combine the answer1 to question1 to answer the combined questions. This means that, to answer the combined question, a responder must first deduce the part of the combined question that refers to the answer1, and then proceed to answer the combined question based on that answer. |
- You cannot include the answer to question 1 in the combined question. |
@@
1. Plain text processing: |
- Accurately recognize all text content in the PDF image without guessing or inferring. |
- Precisely recognize all text in the PDF image without making assumptions in the Markdown format. |
- Maintain the original document structure, including headings, paragraphs, lists, etc. |
2. Formula Processing: |
- Convert all formulas to LaTeX. |
- Enclose inline formulas with $ $. For example: This is an inline formula $ E = mc2̂ $. |
@@
3. Table Processing: |
- Convert all tables to LaTeX format. |
@@
- Convert all Charts to LaTeX format. |
@@
- Ignore figures from the PDF image; do not describe or convert images. |
5. Output Format: |
- Ensure the Markdown output has a clear structure with appropriate line breaks. |
- Maintain the original layout and format as closely as possible. |
@@
1. Question type check |
@@
2. Evidence relevance Check |

### v4

@@ 2.3 Benchmark and evaluation of Retrieval-Augmented Generation
| - Domain | |
| - Total Pages | |
| - Avg.Tokens | |
| - Avg.Data Type | |
@@
| - Text | |
| - Table | |
| - Formula | |
| - Chart | |
| - Reading Order | |
@@
| - String | |
| - Numeric | |
| - Yes/No | |
| - List | |
@@
| - Understanding | |
| - Reasoning | |
| - Single-page | |
| - Multi-page | |
@@ IV.6 Details in different domains
| 1. Question should represent a plausible inquiry that a person (who has not seen the page) might ask about the information uniquely presented on this page. The questions should not reference this specific page directly (by page number, pointing to a specific paragraph or figure, and never refer to the document using phrases like ’in the document’), nor should they quote the text verbatim. They should use natural language reflecting how someone might inquire about the page’s content without direct access. |
| 2. Question must contain all information and context/background necessary to answer without the document. Do not include phrases like "according to the document" in the question. |
| 3. Question must not contain any ambiguous references, such as ’he’, ’she’, ’it’, ’the report’, ’the paper’, and ’the document’. You MUST use their complete names. |
@@
| - The question MUST be detailed and be based explicitly on information in the document. |
| - The question MUST include at least one entity. |
| - The context sentence the question is based on MUST include the name of the entity. For example, an unacceptable context is "He won a bronze medal in the 4 × 100 m relay". An acceptable context is "Nils Sandström was a Swedish sprinter who competed at the 1920 Summer Olympics." |
| - The answer form should be as diverse as possible, including [Yes/No, Numeric, String, List]. |
| - {additional_task_criteria} |
@@
| - The answer to the combined question should be the same as the answer2. |
| - It must combine the answer1 to question1 to answer the combined questions. This means that, to answer the combined question, a responder must first deduce the part of the combined question that refers to the answer1, and then proceed to answer the combined question based on that answer. |
| - You cannot include the answer to question 1 in the combined question. |
@@
| 1. Plain text processing: |
| - Accurately recognize all text content in the PDF image without guessing or inferring. |
| - Precisely recognize all text in the PDF image without making assumptions in the Markdown format. |
| - Maintain the original document structure, including headings, paragraphs, lists, etc. |
| 2. Formula Processing: |
| - Convert all formulas to LaTeX. |
| - Enclose inline formulas with $ $. For example: This is an inline formula $ E = mc2̂ $. |
@@
| 3. Table Processing: |
| - Convert all tables to LaTeX format. |
@@
| - Convert all Charts to LaTeX format. |
@@
| - Ignore figures from the PDF image; do not describe or convert images. |
| 5. Output Format: |
| - Ensure the Markdown output has a clear structure with appropriate line breaks. |
| - Maintain the original layout and format as closely as possible. |
@@
| 1. Question type check |
@@
| 2. Evidence relevance Check |